OpenStudy (mony01):

Is the answer 0 for this problem? Sketch region enclosed by the given curves and find its area. y=cosx, y=2-cosx, 0

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dcosx%3B+y%3D2-cosx Here is a graph of the two

OpenStudy (mony01):

how can i find its area?

OpenStudy (anonymous):

Are you taking trig or calculus?

OpenStudy (mony01):

calculus

OpenStudy (anonymous):

You must find the definite integral of the function cos(x) first. What is the evaluation of the following? $$\int_0^{2\pi} \cos x \, \mathrm{d} x$$

OpenStudy (mony01):

@bloopman is it 0?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Now, once you've graphed the functions, you see you just have to subtract the area under the curve of cos(x) from 2 - cos(x); however, since the indef. integral over cos(x) is 0, we just take the definite integral of 2-cos(x) on the boundary [0, 2pi]. That is: $$\int_0^{2\pi} 2 - \cos x \, \mathrm{d} x$$

OpenStudy (mony01):

i got 0 is it right?